Solve each of the following systems by substitution.
8x – 4y = 16
Y = 2x – 4

Answer by fractalier(6550)

We take the value for y in the second equation
y = 2x – 4 and plug it in to the first equation
8x – 4y = 16 to get
8x - 4(2x – 4) = 16. Simplifying, we get
8x - 8x + 16 = 16 and 16 = 16. Since this is always true, for any value of x, we know that this is a dependent system with an infinite number of solutions.